# TXL Fusion: A Hybrid Machine Learning Framework Integrating Chemical Heuristics and Large Language Models for Topological Materials Discovery

### v1

@@ Abstract
Topological materials, encompassing topological insulators (TIs)hasan2010topological, qi2011topological and topological semimetals (TSMs),liu2014discovery, xu2015discovery represent unconventional quantum phases of matter characterized by nontrivial electronic band topology. Their robust boundary states, protected against perturbations such as disorder or symmetry breaking, give rise to exotic phenomena including the quantum spin Hall effect,kane2005quantum unusual transport properties,qi2011topological and magnetoelectric responseshu2019transport. These unique properties position topological materials as promising candidates for next-generation quantum and spintronic technologies. Since the emergence of the field, a central challenge has been the reliable identification and classification of such materials. Early efforts relied heavily on first-principles calculations combined with topological band theory,xiao2021first, bansil2016band a computationally intensive but powerful route for establishing topological order. The advent of symmetry indicatorspo2017symmetry and topological quantum chemistrybradlyn2017topological represented a major milestone, enabling efficient diagnosis of many topological phases directly from symmetry representations of electronic states. These symmetry-based approaches facilitated high-throughput computational searches, resulting in extensive databases of candidate topological materials and accelerating both theoretical and experimental exploration.vergniory2019complete, zhang2019catalogue, tang2019comprehensive
@@
Over the past decade, machine learning (ML) has become a scalable alternative to symmetry-based approaches for classifying topological materials.schleder2021machine, hong2025discovery, tyner2024machine, zhang2018machine, choudhary2021high Models such as gradient-boosted trees trained on space group (SG), electron count, and orbital-resolved valence descriptors have achieved strong performance,claussen2020detection and neural networks applied to computed XANES spectra have further expanded predictive capabilities.nina2022machine Despite these advances, conventional ML models operate solely on structured numerical inputs, limiting their ability to incorporate unstructured information—such as material descriptions, experimental annotations, or insights from scientific literature. To overcome data and scalability constraints, composition-based heuristics have been proposed, most notably the topogivity score $g(M)$,ma2023topogivity with subsequent extensions such as the inclusion of Hubbard $U$ parameter for magnetic systems.xu2024discovering While efficient and interpretable, these composition-only rules remain insensitive to essential physical features and often struggle to distinguish closely related phases, particularly TSMs and TIs.
@@ 2 TXL Fusion architecture
The pipeline begins with a composition-based heuristic module, adapted from Ma et al.ma2023topogivity, which assigns elemental contribution scores to estimate the likelihood of a material being trivial, TSM, or a TI. This heuristic captures global compositional trends consistent with chemical intuition—where lighter, nonmetallic elements favor trivial phases, while heavier elements such as Bi, Sb, and Te correlate with topological behavior. However, since heuristic trends alone cannot distinguish between TIs and TSMs (see Table 1 in the "Results and Discussion" section)", we complement it by numerical descriptor module.
The numerical descriptor module encodes physically meaningful quantities—such as space group (SG) symmetry, total and parity-resolved electron counts, orbital occupancies, electronegativity differences, and compositional ratios—into a fixed-length vector. These descriptors, selected through the feature analysis described in Section "Dataset and Feature Selection" and SI Section S1, provide a systematic and interpretable representation of the material’s underlying physics.
@@ 3 Results and Discussion
In this section, we assess the capability of the proposed TXL Fusion model to distinguish topological materials from trivial ones and benchmark it against two standalone baselines: the composition-based heuristic rule $g(M)$ and a numerical descriptor-based XGB model. Detailed training procedures are provided in the Supporting Information (Section S3). The dataset was divided into 80% for training and 20% for testing, with the latter designated as Discovery Space–1, on which high-throughput screening was conducted using features learned from the training data. The training subset was further partitioned into 80% sub-training and 20% validation data; no validation results are reported for $g(M)$ since it represents a fixed analytical rule. Model performance is evaluated using precision, recall, and F1-score:
@@
Table 1 summarizes comparative performance. The $g(M)$ rule, while interpretable and chemically motivated, performs poorly in Discovery Space-1: although $g(M)^{\mathrm{Trivial-Others}}$ distinguishes trivial compounds reasonably well (F1 = 0.81), it performs worse on TSMs (F1 = 0.62) and fails entirely for TIs (F1 = 0.00), reflecting its reliance on composition vectors alone. The numerical descriptor based XGB model, leveraging richer numerical and structural features, outperforms $g(M)$ rule for TI and TSM classification (F1 > 0.85 across splits) but still struggles with TIs (F1 = 0.55 in Discovery Space-1), limited by data imbalance and feature sparsity.
By contrast, TXL Fusion consistently outperforms the baselines across all classes, achieving F1 scores of 0.89 (trivial), 0.89 (TSM), and 0.62 (TI). This represents a clear improvement over XGB, particularly for TIs (+0.07). Its superior performance stems from its hybrid design, which integrates complementary information sources to enhance both classification accuracy and generalization.
It is worth mentioning that performance varies with chemical complexity (Supporting Information Section S4). For one-element compounds, both XGB and TXL struggle to capture topological phases (F1 < 0.25). Accuracy improves for 2–3 element compounds, where TXL slightly outperforms XGB (F1 $\approx$ 0.62–0.64 vs. 0.58–0.60). The largest advantage appears for 4-element systems, with TXL sustaining reliable TI predictions (F1 = 0.57) while XGB collapses (F1 = 0.23). For 5–6 element compounds, both models perform well on trivial and TSM classes, but TXL remains more balanced for scarce TIs (F1 = 0.61 and 0.33 vs. XGB’s 0.33 and 0.00).
@@
To further evaluate the TXL Fusion’s predictive capability, we used the set of 1,433 materials reported in Ma et al.ma2023topogivity as a discovery space. These materials, originally curated by Tang et al.tang2019comprehensive, are of particular interest because their topological character cannot be resolved using symmetry indicators. Among them, 1,235 compounds are already present in the topological materials database, leaving 198 as a new testing set, which we refer to as Discovery Space–2. Two compounds belonging to SG 178 were excluded, as this SG is not represented in the topological materials database, resulting in 196 candidate materials.
@@ 4 Concluding remarks
The TXL Fusion framework demonstrates that combining chemically informed heuristics with LLM embeddings yields robust and generalizable predictions of topological character. Across all classes, it consistently outperforms both the standalone XGB and heuristic $g(M)$ baselines, underscoring the effectiveness of hybrid architectures that unify human-crafted rules with data-driven representations. The balanced integration of these complementary components enables the model to capture interpretable chemical trends alongside subtle, high-dimensional patterns within the data. To promote broader accessibility, the TXL Fusion model will soon be available for public use through the Aitomistic Hub at aitomistic.xyz.
Despite these advances, several limitations temper the current performance. Topological insulators remain the most difficult class to predict accurately, a consequence of both intrinsic and extrinsic factors. First, TIs are underrepresented relative to other classes, leading to class imbalance during training. Second, the dataset spans a broad chemical and crystallographic space, with many space groups and element combinations represented by only a handful of examples. As a result, individual TIs often lack closely related analogues from which the model could learn consistent patterns. Compounding this, the DFT-based labels on which the training relies are not always reliable—especially for small-gap or borderline cases—introducing label noise that can render some systems effectively unpredictable. These challenges are reflected in the element-wise breakdown of model performance: compounds with fewer constituent elements (particularly binaries and pure elements) exhibit lower F1-scores for TIs, whereas more compositionally complex materials show markedly better recall and precision for the trivial and semimetal classes.
At the same time, our results reveal important strengths. The model performs best for compounds with intermediate chemical complexity (three to five elements), where structural and electronic descriptors are more distinctive and informative. For these systems, TXL Fusion achieves high accuracy for all classes, suggesting that it effectively leverages richer descriptor space when available. Encouragingly, large-gap TIs—those of greatest experimental interest—tend to be recognized more reliably, underscoring the model’s potential for guiding real-world discovery rather than merely reproducing DFT labels.
Looking forward, several avenues could further improve performance and broaden applicability. Addressing class imbalance through data augmentation or cost-sensitive learning, refining label quality through cross-validation with experimental data, and incorporating transfer learning across chemical families are all promising directions. More broadly, the TXL Fusion paradigm—fusing heuristic chemical knowledge with LLM-driven embeddings—offers a flexible foundation that could be extended beyond topological materials to accelerate the discovery of other quantum or functional materials where data scarcity and complexity remain major bottlenecks.
@@ 5 Code and Data Availability
Comprehensive methodology and supplementary analysis–including feature analysis, TXL Fusion design, training details, element-count performance breakdowns, and DFT computational settings — are available in the Supporting Information (Sections S1–S5).

### v2

@@ Abstract
Topological materials, encompassing topological insulators (TIs)hasan2010topological, qi2011topological and topological semimetals (TSMs),liu2014discovery, xu2015discovery represent unconventional quantum phases of matter characterized by nontrivial electronic band topology. Their robust boundary states, protected against perturbations such as disorder or symmetry breaking, give rise to exotic phenomena including the quantum spin Hall effect,kane2005quantum unusual transport properties,qi2011topological and magnetoelectric responseshu2019transport. These unique properties position topological materials as promising candidates for next-generation quantum and spintronic technologies. Since the emergence of the field, a central challenge has been the reliable identification and classification of such materials. Early efforts relied heavily on first-principles calculations combined with topological band theory,xiao2021first, bansil2016band a computationally intensive but powerful route for establishing topological order. The advent of symmetry indicatorsslager2013space, kruthoff2017topological, po2017symmetry and topological quantum chemistrybradlyn2017topological represented a major milestone, enabling efficient diagnosis of many topological phases directly from symmetry representations of electronic states. These symmetry-based approaches facilitated high-throughput computational searches, resulting in extensive databases of candidate topological materials and accelerating both theoretical and experimental exploration.vergniory2019complete, zhang2019catalogue, tang2019comprehensive
@@
Over the past decade, machine learning (ML) has become a scalable alternative to symmetry-based approaches for classification and properties prediction of topological materials.zhang2018machine, scheurer2020unsupervised, cao2020artificial, choudhary2021high, schleder2021machine, tyner2024machine, hong2025discovery, haosheng2025predicting Models such as gradient-boosted trees trained on space group (SG), electron count, and orbital-resolved valence descriptors have achieved strong performance,claussen2020detection and neural networks applied to computed XANES spectra have further expanded predictive capabilities.nina2022machine Despite these advances, conventional ML models operate solely on structured numerical inputs, limiting their ability to incorporate unstructured information—such as material descriptions, experimental annotations, or insights from scientific literature. To overcome data and scalability constraints, composition-based heuristics have been proposed, most notably the topogivity score $g(M)$,ma2023topogivity with subsequent extensions such as the integration of Hubbard $U$ parameter for magnetic systemsxu2024discovering and the inclusion of quantum correlations across elementsxu2025quantum. While efficient and interpretable, these composition-only rules remain insensitive to essential physical features and often struggle to distinguish closely related phases, particularly TSMs and TIs.
@@ 2 TXL Fusion architecture
The pipeline begins with a composition-based heuristic module, adapted from Ma et al.ma2023topogivity, which assigns elemental contribution scores to estimate the likelihood of a material being trivial, TSM, or a TI. This heuristic captures global compositional trends consistent with chemical intuition—where lighter, nonmetallic elements favor trivial phases, while heavier elements such as Bi, Sb, and Te correlate with topological behavior. However, since heuristic trends alone cannot distinguish between TIs and TSMs (see Table 1 in the "Results and Discussion" section), we complement it by numerical descriptor module.
The numerical descriptor module encodes physically meaningful quantities—such as SG symmetry, total and parity-resolved electron counts, orbital occupancies, electronegativity differences, and compositional ratios—into a fixed-length vector. These descriptors, selected through the feature analysis described in Section "Dataset and Feature Selection" and Supporting Information (Section S1), provide a systematic and interpretable representation of the material’s underlying physics.
@@ 3 Results and Discussion
In this section, we assess the capability of the proposed TXL Fusion model to distinguish topological materials from trivial ones and benchmark it against two standalone baselines: the composition-based heuristic rule $g(M)$ and a numerical descriptor-based XGB model. Detailed training procedures are provided in the Supporting Information (Section S3). The dataset was divided into 80% for training and 20% for testing (7637 materials), with the latter designated as Discovery Space-1, on which high-throughput screening was conducted using features learned from the training data. The training subset was further partitioned into 80% sub-training and 20% validation data; no validation results are reported for $g(M)$ since it represents a fixed analytical rule. Model performance is evaluated using precision, recall, and F1-score:
@@
Table 1 summarizes comparative performance on 7637 unseen materials of Discovery Space-1. The $g(M)$ rule, while interpretable and chemically motivated, performs poorly in Discovery Space-1: although $g(M)^{\mathrm{Trivial-Others}}$ distinguishes trivial compounds reasonably well (F1 = 0.81), it performs worse on TSMs (F1 = 0.62) and fails entirely for TIs (F1 = 0.00), reflecting its reliance on composition vectors alone. The numerical descriptor based XGB model, leveraging richer numerical and structural features, outperforms $g(M)$ rule for TI and TSM classification (F1 > 0.85 across splits) but still struggles with TIs (F1 = 0.57 in Discovery Space-1), limited by its factorized nature, which treats features such as SG, electron count, orbital occupations, and elemental ratios as independent scalars.
By contrast, TXL Fusion consistently outperforms the baselines across all classes, achieving F1 scores of 0.89 (trivial), 0.89 (TSM), and 0.62 (TI). This represents a clear improvement over XGB, particularly for TIs by +5%. This superior performance stems from its hybrid design, with its primary advantage arising from LLM embeddings that capture complex, coupled electronic–structural correlations, allowing the model to surpass the limitations of factorized numerical descriptors. A detailed comparison of semantic versus numerical descriptors, highlighting these advantages, is provided in Section S4 of the Supporting Information.
It is worth mentioning that performance varies with chemical complexity (Supporting Information Section S5). For single-element compounds, both TXL and XGB struggle with TIs (F1 < 0.25) due to data scarcity, limited descriptor variability, and subtle, non-compositional physical effects. Accuracy improves for 2–3 element compounds, where TXL slightly outperforms XGB (F1 $\approx$ 0.62–0.64 vs. 0.59–0.63). The largest advantage appears for 4–5 element systems, where TXL sustains reliable TI predictions (F1 = 0.57–0.61) while XGB sharply underperforms (F1 = 0.24–0.31). For six-element compounds—the most sparsely represented category in the dataset—TXL Fusion retains better balance for rare TI cases (F1 = 0.33 compared to 0.00 for XGB), while both models achieve strong performance for trivial and TSM classes.
@@
To further evaluate the TXL Fusion’s predictive capability, we used the set of 1,433 materials reported in Ma et al.ma2023topogivity as a discovery space. These materials, originally curated by Tang et al.tang2019comprehensive, are of particular interest because their topological character cannot be resolved using symmetry indicators. Among them, 1,235 compounds are already present in the topological materials database, leaving 198 as a new testing set, which we refer to as Discovery Space-2. Two compounds belonging to SG 178 were excluded, as this SG is not represented in the topological materials database, resulting in 196 candidate materials.
@@ 4 Concluding remarks
In this work, we introduced TXL Fusion, a hybrid framework that integrates chemically informed heuristics, numerical descriptors, and LLM-derived semantic embeddings for robust classification of topological materials. Across all classes, TXL Fusion consistently surpasses both the standalone XGB model and the heuristic $g(M)$ rule, demonstrating the effectiveness of combining domain intuition with data-driven, context-aware representations. By unifying interpretable chemical trends with high-dimensional semantic embeddings, the model achieves improved accuracy and generalization, particularly for compositionally complex systems. To promote broader community access, TXL Fusion will be made publicly available through the Aitomistic Hub at aitomistic.xyz.
Despite these advances, several intrinsic challenges remain. TIs are the most difficult class to predict, primarily due to label uncertainty and data sparsity. As discussed in Section S7 of the Supporting Information, DFT-based topological labels—especially for small-gap or borderline systems—can be unreliable, effectively introducing noise into the training data. Consistent with this interpretation, evaluation on high-confidence TI subsets shows systematic performance improvements as label reliability increases (Fig. S7). In addition, element-resolved analysis in Section S5 reveals reduced performance for low-complexity compounds (single- and binary-element systems), arising from a combination of data scarcity, limited descriptor expressiveness due to compositional features collapsing to constants or narrow ranges, and the dominance of subtle, non-compositional physical effects (see Section S5.1).
While these limitations could be mitigated by incorporating more localized or symmetry-resolved physical descriptors—such as symmetry-indicator eigenvalues, band inversion indicators, Wannier charge centers, or orbital-projected band-edge features—such quantities are computationally intensive and not uniformly available at scale. Within our framework, we therefore prioritize broad chemical coverage, enabling large-scale screening based solely on elemental information without requiring prior expensive calculations.
At the same time, TXL Fusion demonstrates strong generalization under chemical sparsity. As detailed in Section S5.2, the model maintains reliable performance for trivial and TSM phases even in sparsely sampled or partially out-of-distribution chemical families, with its advantage over numerical baselines increasing alongside chemical and electronic complexity. Moreover, confidence-based calibration analysis (Section S5) shows that TXL Fusion is well calibrated at high confidence levels, indicating that its most confident predictions are highly reliable and well suited for guiding experimental validation.
More broadly, this study highlights the power of LLM-based semantic embeddings for materials discovery. As discussed in Section S4, these embeddings capture coupled electronic–structural relationships that are fragmented in conventional factorized numerical descriptors, enabling improved discrimination of subtle topological regimes. Looking forward, continued improvements in label fidelity, expanded high-confidence datasets, and the incorporation of additional physics-aware learning strategies—such as transfer learning across chemical families or cost-sensitive training to address class imbalance—are expected to further enhance performance. More generally, the TXL Fusion framework is not limited to topological materials and provides a flexible foundation for accelerating discovery across a broad range of quantum and functional materials systems.
@@ 5 Code and Data Availability
Comprehensive methodology and supplementary analyses—including feature engineering, TXL Fusion architecture and training details, a semantic versus numerical descriptor performance comparison, element-resolved performance and confidence analysis with dedicated discussions of low-complexity compounds and chemical sparsity generalization, DFT computational settings, and an assessment of noise and reliability in DFT-based topological labels—are provided in the Supporting Information (Sections S1–S7).